What are three consecutive integers such as that the square of the third is 5 more than 4 times the sum of the first and second?
Answer by CubeyThePenguin(3113)

x-1, x, x+1


(x+1)^2 = 5 + 4(x-1 + x)
x^2 + 2x + 1 = 5 + 8x - 4
x^2 - 6x = 0
x(x - 6) = 0
x = 0, x = 6

The integers could be {-1, 0, 1} or {5, 6, 7}.